OpenStudy (anonymous):

Classify an equiangular triangle by its angles. Select one: a. obtuse b. cannot be determined c. right d. acute

OpenStudy (anonymous):

Image?

OpenStudy (anonymous):

@AnonymousBeast

OpenStudy (anonymous):

All that gives me is a login page. Use the Draw tool to model it as well as you can.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A right triangle has exactly one angle measure of 90 degrees An obtuse triangle has exactly one angle measure greater than 90 degrees An acute triangle has all angles less than 90 degrees. We know that 58 and 79 are acute angles, but we need to find the last. We know that all triangles have interior angle sums of 180 degrees. So: 58+79+x=180 137+x=180 -137 -137 x=43 degrees Therefore, the angle measures are 58,79,43. It is an acute triangle.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Anytime.